# Revisiting In-context Learning Inference Circuit in Large Language Models

### v1

@@ 4.2 Step 3, Induction Head: Feature Retrieval on Task Subspace
Induction is Correct in Minority Subspaces. Similar to Fig. 5 (Middle), we mark (1) the attention heads with the sum of attention scores from the query’s forerunner token $s_{q}$ (as attention query) to all the label tokens $[y_{1}],\dots,[y_{k}]$ (as attention keys) in the demonstration more than $5k/n_{t}$ as induction heads, and (2) attention heads with the sum of scores to all the correct label tokens more than $5k/|\mathbb{Y}|n_{t}$ as correct induction heads ($\mathbb{Y}$ is a label space). We show the number of both kinds of induction heads in Fig. 6 (Left, detailed head statistics in Appendix F.1), where a unimodal pattern is observed later than the copy processing of Step 2. Moreover, more than half of the induction heads are not correct ones, suggesting that task-specific feature similarity can only be caught on some induction subspaces (defined by low-rank transition matrix $W_{Q}^{h\top}W_{K}^{h}$ of correct induction head $h$). We enhance this claim in Fig. 6 (Middle) (details in Appendix A.4), where both vanilla attention (without transformation and head split) and attention scores averaged among heads show low assignment on correct label tokens, while some heads show considerable correctness. Considering the average value, the majority of attention heads almost randomly copy label token information to the query, causing the prediction biased to the frequency of labels in the prompt (Zhao et al., 2021). As the reason, we infer that the hidden states are sufficient (Fig. 3) but not minimum for ICL, where redundant information interferes with the similarity calculation of attention.
@@
Figure 7: Label representations of $k$ demonstrations visualized on (Left 4) correct and (Right 4) wrong induction head, on one sample of SST-2 (see Appendix F.3). $\circ$: “positive” label, $\times$: “negative” label, $\blacktriangle$: zero vector. Color: attention assigned to query, negative to positive (cartography: Appendix A.5).
@@
The results are shown in Fig. 6 (Right), where: (1) A significant overlap of induction heads indicates that a part of correct induction heads is inherent in the model, built by the pre-training process (Reddy, 2024; Singh et al., 2024b). (2) Such overlap is not fully observed, suggesting that some induction subspaces are task-specific: input texts evoke task-specific attention in induction heads, enabling the anisotropy multiplex of different sub-spaces in the hidden spaces to transmit relevant information for various tasks. Therefore, we can restore ICL to implicit end-to-end multi-task learning methods with hidden state multiplexing since they also use various task heads on common bottom network layers and informative hidden state (Zhang & Yang, 2021).
Demonstration Saturates on Induction Subspace. We visualize the demonstrations’ label token representations mapped on the induction subspaces by the transition matrix $W_{Q}^{h\top}W_{K}^{h}$ and principal component analysis in Fig. 7, indicating: (1) Compared to a correct induction head, a wrong induction head is easier to map label representations linearly inseparably. (2) In the early stage of demonstration ($k=1\rightarrow 2$), when a new demonstration is given, the morphology of attention assignment towards query changes significantly (shown as background color in Fig. 7; see §A.5 for details), while in the late stage ($k=15\rightarrow 16$), attention assignment morphology is stable. This can explain the demonstration saturation (Agarwal et al., 2024; Bertsch et al., 2024): the performance is submodular against the demonstrations. Intuitively, since demonstrations follow a prior distribution, representation of a new demonstration is likely to be located within the closure of existing demonstrations, making it less contributable to the attention assignment in the induction subspace.
@@ 5.1 Ablation Analysis
Table 1: Accuracy variation (%) with each inference step ablated on Llama 3 8B. Small numbers are controlled results of randomly ablating equivalent amounts of connections. Ablations are applied from the bottom to the top layers, and results with various ablated layers are reported.
@@
– Step1: Summarize –
@@ 5 Label $y_{i}$rightarrow Query Forerunner $s_{q}$
To demonstrate that our 3-phase circuit dominates or at least participates in ICL process, we disconnect the related attention connection of each step in the proposed circuit, and test the accuracies without such connections as shown in Table 1. The results show that when the non-trivial connections designated by the proposed circuit are ablated, accuracies of ICL significantly decrease, supporting the existence of our circuit. However, the result doesn’t fully match expectations, for example, the result without induction (line 5) should be consistent with zero-shot (line 6), since all the expected communication from demonstration to query is intercepted, but that’s not the case; and the contribution of Step 1 in later layers are unexpectedly high, indicating the existence of some bypass mechanisms parallelly contributing to ICL accuracies.
@@ 5.2 Bypass Mechanism
Direct Decoding. Residual connection to output embedding allows intermediate hidden states to be decoded directly. Intuitively, a shortcut from Step 1 encodings to the LM head enables ICL with zero-shot capacities, since we have confirmed that encoded representations are informative for ICL tasks, while the decoding methods should be selected carefully (Cho et al., 2024) (e.g. with essential calibration). On the other hand, shortcuts from insufficiently encoded features may lead to meaningless information decoded by language model heads, causing prediction bias, i.e., even if no query is given, ICL still returns unbalanced results (Zhao et al., 2021) decoded from tokens of prompt template (see Appendix D).
@@
Shortcut Induction. Note that a $k$-shot ICL input sequence always contains a $(k-1)$-shot sequence, where the $k$-th forerunner (once served as the query) is previously processed. So every forerunner can directly retrieve previously processed forerunners of demonstrations to copy their induction results directly. The non-trivial result (Table 2) with all forerunners disconnected from each other confirms such infer.
@@ 5.3 Explaination towards Observed ICL Phenomena
Difficulty-based Demonstration Selection. In §3.1, we find that in the zero-shot scenario, perplexed texts are harder to be encoded, which explains the observation of PPL-ICL (Gonen et al., 2023) selecting demonstrations with lower perplexity. Moreover, while the demonstrations increase, LMs can encode more complex inputs with diverse information to update the attention assignment shown in Fig. 7, making it beneficial to input harder demonstrations later, which explains the ICCL (Liu et al., 2024), which build demonstrations sequence from easy to hard.
Prediction Bias. (1) Contextual Bias: As shown in §5.2 and Appendix D, direct decoding insufficiently encoded information adds meanless logits into LM’s output, causing a background prediction value even if no queries are given (named bias). (2) Position Bias: As shown in §3.2, closer input texts are encoded more similarly, so label tokens near the query have more similar information to the query, causing more attention assignment in the induction processing, so that more influences on the prediction. (3) Frequency Bias: As shown in §4.2, in the induction, some attention heads are without selectivity towards labels, causing an averaged induction from label tokens to the query, triggering a prediction bias towards the label frequency in the demonstration, even if their contribution (absolute value of attention score on label tokens) is small. All three biases are observed by Zhao et al. (2021), and can be removed by ICL calibration methods.
@@ 6 Conclusion and Discussion
Inference Acc. Para.# Speed
@@
The Role of Early and Later Layers. Our framework and Fig. 3 show: encoding result of Step 1 can be directly used for classification with reliable decoding, and later transformer layers are not contributing to centroid classification accuracies, leading to a taxonomy of Encoding for Step 1 and Output Preparation for Step 2 and 3. LMs complete multi-task classification implicitly in early layers, and verbalize it by merging task-specific label semantics in later layers. Therefore, we suggest removing some top layers and using a centroid classifier (Cho et al., 2024) to accelerate ICL inference as shown in Table 3 and Appendix D.
@@
Pre-training Possibility from Natural Language Data. A large gap can be considered between such a precise circuit and gradient descent pre-training on the wild data. However, we believe the wild training target contains the ICL circuit functionally. Based on the previous works finding trainability of ICL on linear representation-label pairs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b), we speculate that in early training step, Transformers learn to extract linear representations shown in §3 from wild data (Appendix B), serving as the training input of later layers to evoke the emergence of induction heads with the same mechanism shown in aforementioned previous works. Moreover, our conclusion of Step 3 highlights the input data requirements for the later layers: These data should activate the multiplex of hidden space, i.e., it should implicate multi-task classification with a wide distribution, which is consistent with the aforementioned previous works.
Limitations. (1) These 3 basic operations are not functionally indivisible. Ideally, one can reduce every operation in ICL inference to the interconnection of special attention heads to ulteriorly examine how the operating subspaces interact between steps. (2) Even though we have some insights into how pre-training promotes ICL in natural language models, its detailed dynamics are still unknown. One can start by decomposing pre-training targets into implicit tasks, and examine how these tasks can evoke the occurrence of the 3-step inference operations. (3) We only focus on classification tasks, while we believe that our findings can be applied to non-classification tasks, efforts are still needed to fill the gap.
@@ Appendix B LM Pre-training Dynamics Measured by ICL Circuit
One can divide a self-regression model into an early part and a later part, where the early part encodes the input into a hidden representation, and the later part decodes the hidden representation back to the input. So, the training object can also be divided into an encoding loss and a decoding loss. According to the discussion in §6, the summarization operation (Step 1) can be classified as encoding, and the other two steps of the induction circuit can be classified as decoding. Intuitively, since the decoding operations require the encoding results as input, unless the encoding operation converges to a stable output, the decoding can not be trained since the input-output mapping is noised, causing unstable gradients to interfere with the training (Liu et al., 2020).
@@ B.1 Data Distribution Requirement Explained by Hidden State Multiplex
Figure 11: The 3 operating magnitudes on Pythia 6.9B with various pre-training steps. Left: Step 1, summarize; Middle: Step 2, semantics merge; Right: Step 3, feature retrieval and copy, measured by the correct induction head numbers.
@@
During the training, (1) When the label space is small, the trained attention heads only need to extract the projected length of the query on each label eigen-subspace. For each label, such operation has a parameters’ analytical solution with (encoding kernel) $W_{Q}^{\top}W_{K}=I$ and (decoding transformation) $W_{O}W_{V}=o_{y}^{\top}e_{y}$, where $o_{y}$ is the label token’s output embedding, and $e_{y}$ is the label’s eigen-subspace. From such an operation, theoretically, one layer can handle at most $|\mathcal{H}|$ labels, where $|\mathcal{H}|$ is the head amounts. While, considering the sparsity of these eigen-subspaces, such an upper bound can be increased to $d^{\prime}|\mathcal{H}|$ by multiplex one head to decode an orthogonal group of labels with orthogonal eigen-subspaces of $\mathcal{E}=[e_{y_{1}};e_{y_{2}};\dots;e_{y_{d^{\prime}}}]$ and orthogonal output embedding $\mathcal{O}=[o_{y_{1}};o_{y_{2}};\dots;o_{y_{d^{\prime}}}]$, where $d^{\prime}$ is the inner dimension of attention head, with decoding transformation $W_{O}W_{V}=\mathcal{O}^{\top}\mathcal{E}$. (2) When the label space expands [^10] Notice that such a situation can also occur when the orthogonality between eigen-subspaces or output embeddings is lost. A common situation is that the variance of the cluster increases, creating confusion within the decoding space of the attention head., the decoding transformation can not distinguish all the clusters since the $W_{O}W_{V}$ is low-rank. Driven by the training loss, the model can choose to transform the encoding kernel to focus on catching the most similar label tokens with the query, and copy the label tokens’ information back to the query. As a result, one attention head can catch at most $d^{\prime}$ groups of label tokens mapped collinearly by the encoding kernel (note that this set of labels may not appear simultaneously in the context, so confusion can be avoided), and the common space of these label words become the induction subspace shown in §4.2.
@@
Notice that this section is our hypothesis based on the results of this paper, which should be empirically validated in future work. In addition, we still need to find implicit input-label tuples in wild data, which will also be addressed in future work.
Appendix C Can LMs Segment ICL Prompt?
@@
As a preliminary observation, we visualize the attention scores with the last forerunner token as the attention queue at layer 21 (a layer with high encoding magnitude, refer to Fig. 2) from an input case, as shown in Fig. 13 (Left), where most of the attention heads are focused on the query text. The visualization suggests that encoding operations are localized into the query.
[…]
@@
Appendix D Measurement on Direct Decoding

### v2

@@ 4.2 Step 3, Induction Head: Feature Retrieval on Task Subspace
Induction is Correct in Minority Subspaces. Similar to Fig. 5 (Middle), we mark (1) the attention heads with the sum of attention scores from all the label tokens $[y_{1}],\dots,[y_{k}]$ (as attention keys) to the query’s forerunner token $s_{q}$ (as attention query) in the demonstration more than $5k/n_{t}$ as induction heads, and (2) attention heads with the sum of scores from all the correct label tokens more than $5k/|\mathbb{Y}|n_{t}$ as correct induction heads ($\mathbb{Y}$ is a label space). We show the number of both kinds of induction heads in Fig. 6 (Left, detailed head statistics in Appendix H.1), where a unimodal pattern is observed later than the copy processing of Step 2. Moreover, more than half of the induction heads are not correct ones, suggesting that task-specific feature similarity can only be caught on some induction subspaces (defined by low-rank transition matrix $W_{Q}^{h\top}W_{K}^{h}$ of correct induction head $h$). We enhance this claim in Fig. 6 (Middle) (details in Appendix A.4), where both vanilla attention (without transformation and head split) and attention scores averaged among heads show low assignment on correct label tokens, while some heads show considerable correctness. Considering the average value, the majority of attention heads almost randomly copy label token information to the query, causing the prediction biased to the frequency of labels in the prompt (Zhao et al., 2021). As the reason, we infer that the hidden states are sufficient (Fig. 3) but not minimum for ICL, where redundant information interferes with the similarity calculation of attention.
@@
Figure 7: Label representations of $k$ demonstrations visualized on Left 4: correct and Right 4: wrong induction head, on one sample of SST-2 (see Appendix H.3). $\circ$: “positive” label, $\times$: “negative” label, $\blacktriangle$: zero vector. Color: attention assigned to query, negative to positive (cartography: Appendix A.5).
@@
The results are shown in Fig. 6 (Right), where: (1) A significant overlap of induction heads indicates that a part of correct induction heads is inherent in the model, built by the pre-training process (Reddy, 2024; Singh et al., 2024b). (2) Such overlap is not fully observed, suggesting that some induction subspaces are task-specific: input texts evoke task-specific attention in induction heads, enabling the anisotropy multiplex of different subspaces in the hidden spaces to transmit relevant information for various tasks. Therefore, we can restore ICL to implicit end-to-end multi-task learning with hidden state multiplexing since they also use various task heads on common bottom network layers and informative hidden states (Zhang & Yang, 2021).
Demonstrations Saturate on Induction Subspace. We visualize the demonstrations’ label token representations mapped on the induction subspaces by the transition matrix $W_{Q}^{h\top}W_{K}^{h}$ and principal component analysis in Fig. 7, indicating: (1) Compared to a correct induction head, a wrong induction head is easier to map label representations linearly inseparably. (2) In the early stage of demonstration ($k=1\rightarrow 2$), when a new demonstration is given, the morphology of attention assignment towards query changes significantly (shown as background color in Fig. 7; see §A.5 for details), while in the late stage ($k=15\rightarrow 16$), attention assignment morphology is stable. This can explain the demonstration saturation (Agarwal et al., 2024; Bertsch et al., 2024): the performance is submodular against the demonstrations. Intuitively, since demonstrations follow a prior distribution, representation of a new demonstration is likely to be located within the closure of existing demonstrations, making it less contributable to the attention assignment in the induction subspace.
@@ 5.1 Ablation Analysis
Table 1: Accuracy variation (%) with each inference step ablated on Llama 3 8B. Small numbers are controlled results ($\mathrm{mean}\pm\mathrm{std}$) of randomly ablating equivalent amounts of connections. Ablations are applied from the bottom to the top layers, and results with various ablated layers are reported (detailed settings: Appendix A.7).
@@
– Step1: Input Text Encode –
@@ 5 Label $y_{i}$rightarrow Query Forerunner $s_{q}$
To demonstrate that our 3-phase circuit dominates or at least participates in ICL process, we disconnect the related attention connection of each step in the proposed circuit (see Appendix A.7 for details), and test the accuracies without such connections as shown in Table 1. The results show that: compared to the controlled results where trivial connections are ablated, when the non-trivial connections designated by the proposed circuit are ablated, accuracies of ICL significantly decrease, supporting the existence of our circuit. However, the result doesn’t fully match expectations, for example, the result without induction (line 5) should be consistent with zero-shot (line 6), since all the expected communication from demonstration to query is intercepted, but that’s not the case; and the contribution of Step 1 in later layers are unexpectedly high, indicating the existence of some bypass mechanisms parallelly contributing to ICL accuracies.
@@ 5.2 Bypass Mechanism
Direct Decoding. Residual connection to output embedding allows intermediate hidden states to be decoded directly. Intuitively, a shortcut from Step 1 encodings to the LM head enables ICL with zero-shot capacities, since we have confirmed that encoded representations are informative for ICL tasks (§3.2), while the decoding methods should be selected carefully (Cho et al., 2024) (e.g. with essential calibration). On the other hand, shortcuts from insufficiently encoded features may lead to meaningless information decoded by language model heads, causing prediction bias, i.e., even if no query is given, ICL still returns unbalanced results (Zhao et al., 2021) decoded from tokens of prompt template (see Appendix E for details).
@@
Shortcut Induction. Note that a $k$-shot ICL input sequence always contains a $(k-1)$-shot sequence, where the $k$-th forerunner (served as the $(k-1)$-shot query) is previously processed. So every forerunner can directly retrieve previously processed forerunners of demonstrations to copy their induction results directly. The non-trivial result (Table 2) with all forerunners disconnected from each other confirms such infer.
@@ 5.3 Explaination towards Observed ICL Phenomena
Difficulty-based Demonstration Selection. In §3.1, we find that in the zero-shot scenario, perplexed texts are harder to encode, explaining the observation of PPL-ICL (Gonen et al., 2023): selecting demonstrations with lower perplexity can improve ICL performance. Moreover, while the demonstrations increase, LMs can encode more complex inputs with diverse information to update the attention assignment shown in Fig. 7, making it beneficial to input harder demonstrations later, which explains the ICCL (Liu et al., 2024), which build demonstrations sequence from easy to hard.
Prediction Bias. (1) Contextual Bias: As shown in §5.2 and Appendix E, direct decoding insufficiently encoded information adds meanless logits into LM’s output, causing a background prediction value even if no queries are given (named bias). (2) Positional Bias: As shown in §3.2, closer input texts are encoded more similarly, so label tokens near the query have more similar information to the query, causing more attention assignment in the induction processing, so that more influences on the prediction. (3) Frequency Bias: As shown in §4.2, in the induction, some attention heads are without selectivity towards labels, causing an averaged copy processing from label tokens to the query, triggering a prediction bias towards the label frequency in the demonstration, even if their contribution (absolute value of attention score on label tokens) is small. All three biases are observed by Zhao et al. (2021), and can be removed by ICL calibration methods.
@@ 6 Conclusion and Discussion
Inference Acc. # Para. Speed
@@
The Role of Early and Later Layers. Our framework and Fig. 3 show: encoding result of Step 1 can be directly used for classification with reliable decoding, and later transformer layers are not contributing to centroid classification accuracies, leading to a taxonomy of Encoding for Step 1 and Output Preparation for Step 2 and 3: LMs complete multi-task classification implicitly in early layers, and verbalize it by merging task-specific label semantics in later layers. Therefore, we suggest an early-exiting inference: removing some top layers and using a centroid classifier (Cho et al., 2024) to accelerate ICL as shown in Table 3 and Appendix E.
@@
Pre-training Possibility from Natural Language Data. A large gap can be considered between such a delicate circuit and gradient descent pre-training on the wild data. However, we believe the wild training target contains the ICL circuit functionally. Based on the previous works finding trainability of ICL on linear representation-label pairs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b), we speculate that in early training step, Transformers learn to extract linear representations shown in §3 from wild data (Appendix B), serving as the training input of later layers to evoke the emergence of induction heads with the same mechanism shown in aforementioned previous works. Moreover, our conclusion of Step 3 highlights the input data requirements for the later layers: these data should activate the multiplex of hidden space, i.e., it should implicate multi-task classification with a wide distribution, which is consistent with the aforementioned previous works.
Comparison with Previous Works. Several prior studies have sought to interpret ICL as known algorithms, including implicit gradient descent (Dai et al., 2023), kernel regression (Han et al., 2023a), and implicit Bayesian inference (Xie et al., 2022), etc. However, as noted in §2.1, these approaches fall short of fully explaining the phenomena observed during the ICL inference process. For instance, gradient descent is known to be fragile against label noise (Zhang et al., 2021), leading to a misalignment when analogies are drawn to ICL, which is robust against label noise (Min et al., 2022). Similarly, attempts to explain ICL as kernel regression fail to account for positional bias, making them disconnected from empirical studies that document various inference phenomena in real-world LMs. Also, other works have employed induction circuits to explain ICL dynamics (Wang et al., 2023; Elhage et al., 2021; Olsson et al., 2022), yet significant gaps remain in aligning these explanations with empirical observations. Our work is the first to unify the fragmented conclusions from prior empirical studies (discussed in §5.3) through detailed experimental measurements. By demonstrating the alignment of these observations, we emphasize the novelty and primary contribution of our paper.
Limitations. (1) These 3 basic operations are not functionally indivisible. Ideally, one can reduce every operation in ICL inference to the interconnection of special attention heads to ulteriorly examine how the operating subspaces interact between steps, and reconstruct ICL behavior from a minimal set of attention heads. Also, although we show the significance of the inference circuit in the ablation analysis (§5.1), measuring the connectivity of these attention heads can also be beneficial to get more insights into the circuit. (2) The conclusions may not align with scenarios where ground-truth labels are not provided in the context in some cases, which are often referred to as in-weight learning, where significant differences or even antagonism with standard ICL have been highlighted by previous works (Chan et al., 2022; Reddy, 2024) [^9] In these works, toy models even struggle to work on both inputs simultaneously, so that it is natural to consider different inference circuit for the in-weight learning., reasonably and necessarily warranting separate discussion (discussed in Appendix G). This paper explains the inference behavior of the model under ICL conditions, leaving the in-weight learning scenario for future works. (3) We only focus on classification tasks, while we believe that our findings can be applied to non-classification tasks, efforts are still needed to fill the gap.
Reproducibility Statement
Author Contributions
Hakaze Cho, also known as Yufeng Zhao, handled the entire workload in this paper. He provided ideas, designed / conducted experiments, collected / described data, and wrote / revised the paper.
N.I., as their supervisor, provided valuable feedback, revisions, and material research support. M.K. and Y.S. actively participated in discussions and offered some helpful suggestions.
Acknowledgments
This work is supported by The Nakajima Foundation Start-Up Support Grant.
Appendices
@@ A.3 Training and Inference of Centroid Classifier
Figure 10: Visualization of non-trivial attention connection defined and disconnected in ablation analysis in §5.1 and Table 1. Notations are same as Table 1.
A.7 Experiment Setting of Ablation Experiments in Table 1
In Table 1, we attribute each step of the inference process to specific attention connections, also shown in Fig. 10. When we aim to remove this step from the inference, we eliminate (i.e. zeroing) all corresponding attention connections from layer 0 to layer $\{25\%,50\%,75\%,100\%\}\times\mathrm{TotalLayers}$.
For reference, for each experiment, we also conduct controlled experiments where the same amount of randomly selected attention connections are removed from the same layers as the experimental values. The controlled results are shown as smaller numbers under the experimental results.
@@ Appendix B LM Pre-training Dynamics Measured by ICL Circuit
One can divide a self-regression model into an early part and a later part, where the early part encodes the input into a hidden representation, and the later part decodes the hidden representation back to the input. So, the training object can also be divided into an encoding loss and a decoding loss. According to the discussion in §6, the operation of Step 1 can be classified as encoding, and the other two steps of the induction circuit can be classified as decoding. Intuitively, since the decoding operations require the encoding results as input, unless the encoding operation converges to a stable output, the decoding can not be trained since the input-output mapping is noised, causing unstable gradients to interfere with the training (Liu et al., 2020).
@@ B.1 Data Distribution Requirement Explained by Hidden State Multiplex
Figure 12: The 3 operating magnitudes on Pythia 6.9B with various pre-training steps. Left: Step 1, Input Text Encode; Middle: Step 2, Semantics Merge; Right: Step 3, Feature Retrieval and Copy, measured by the correct induction head numbers.
@@
During the training, (1) When the label space is small, the trained attention heads only need to extract the projected length of the query on each label’s eigen-subspace. For each label, such operation has a parameters’ analytical solution with (encoding kernel) $W_{Q}^{\top}W_{K}=I$ and (decoding transformation) $W_{O}W_{V}=o_{y}^{\top}e_{y}$, where $o_{y}$ is the label token’s output embedding, and $e_{y}$ is the label’s eigen-subspace. From such an operation, theoretically, one layer can handle at most $|\mathcal{H}|$ labels, where $|\mathcal{H}|$ is the head amounts. While, considering the sparsity of these eigen-subspaces, such an upper bound can be increased to $d^{\prime}|\mathcal{H}|$ by multiplex one head to decode an orthogonal group of labels with orthogonal eigen-subspaces of $\mathcal{E}=[e_{y_{1}};e_{y_{2}};\dots;e_{y_{d^{\prime}}}]$ and orthogonal output embedding $\mathcal{O}=[o_{y_{1}};o_{y_{2}};\dots;o_{y_{d^{\prime}}}]$, where $d^{\prime}$ is the inner dimension of attention head, with decoding transformation $W_{O}W_{V}=\mathcal{O}^{\top}\mathcal{E}$. (2) When the label space expands [^12] Notice that such a situation can also occur when the orthogonality between eigen-subspaces or output embeddings is lost. A common situation is that the variance of the cluster increases, creating confusion within the decoding space of the attention head., the decoding transformation can not distinguish all the clusters since the $W_{O}W_{V}$ is low-rank. Driven by the training loss, the model can choose to transform the encoding kernel to focus on catching the most similar label tokens with the query, and copy the label tokens’ information back to the query. As a result, one attention head can catch at most $d^{\prime}$ groups of label tokens mapped collinearly by the encoding kernel (note that this set of labels may not appear simultaneously in the context, so confusion can be avoided), and the common space of these label words become the induction subspace shown in §4.2.
Notice that this section is our hypothesis based on the results of this paper, the detailed dynamics are still unknown, which should be empirically validated in future work. One can start by decomposing pre-training targets into implicit tasks, and examine how these tasks can evoke the occurrence of the 3-step inference operations. A possible beginning is: finding implicit input-label tuples in wild data.
Appendix C Can LMs Segment ICL Prompt?
@@
As a preliminary observation, we visualize the attention scores with the last forerunner token as the attention queue at layer 21 (a layer with high encoding magnitude, refer to Fig. 2) from an input case, as shown in Fig. 14 (Left), where most of the attention heads are focused on the query text. The visualization suggests that encoding operations are localized into the query tokens.
@@
[…]
@@
Appendix D Semantics Merge is Non-selective on Location
To investigate whether the copy processing described in §4.1 has selectivity on the forerunner token, on every attention head of each layer, for a given token position $i$, we measure the normalized copy magnitude shown below:
$\mathrm{NCM}_{n_{t}}({\mathcal{\alpha}})=n_{t}{\alpha}_{(i-1)\rightarrow i},$ (18)
where the ${\alpha}_{(i-1)\rightarrow i}$ is the attention score with $i$-th token serving as the attention query and $(i-1)$-th token serving as the attention key. For each layer, we export the $\mathrm{NCM}$ at all positions and on all attention heads, and separately statistics the cases where the $i$-th token is a label token or a non-label token. The results for 4 models on SST-2 are shown in Fig. 15.
From the results, no significant statistical differences between these two types of tokens can be observed, suggesting that the semantics copy process, which is identified as Step 2 of our circuit, is not selective on the token types. However, even if we demonstrate that the model cannot exhibit selectivity in the copy positions within the current prompt, a potential direction for future research is to investigate whether any forerunner tokens can enhance or weaken this copy process. Given that this copying is known to be related to label denoising (§4.1), exploring this and carefully designing these forerunner tokens could significantly benefit the control of ICL behavior.
Figure 15: Copy magnitude normalized by the sequence length from the previous token to the non-label token and label token, for every token and head in each layer. Significant statistical differences cannot be observed.
Appendix E Measurement on Direct Decoding